OpenStudy (anonymous):

i need help on this one........anyone want to help?????pleasee!!!:( If you shift the function F(x) = 2x down four units, what is the new function, G(x)?

OpenStudy (anonymous):

There really isn't enough information... Do you have answer choices?

OpenStudy (anonymous):

If you shift the function F(x) = 2x down four units, what is the new function, G(x)? A.G(x) = 2x - 4 B.G(x) = 4 2x C.G(x) = 4 - 2x D.G(x) = 2x - 4

OpenStudy (anonymous):

did u got them?

OpenStudy (anonymous):

I would have to say A.

OpenStudy (anonymous):

u sure?????? is cuz probly today im gunna graduate and i jus need one more test:( can u please help me?????